# Making Differential Privacy Easier to Use for Data Controllers and Data Analysts using a Privacy Risk Indicator and an Escrow-Based Platform

### v2

@@ Abstract.
Differential privacy (DP) enables private data analysis but is hard to use in practice. For data controllers who decide what output to release, choosing the amount of noise to add to the output is a non-trivial task because of the difficulty of interpreting the privacy parameter $\epsilon$. For data analysts who submit queries, it is hard to understand the impact of the noise introduced by DP on their tasks.
To address these two challenges: 1) we define a privacy risk indicator that indicates the impact of choosing $\epsilon$ on individuals’ privacy and use that to design an algorithm to choose $\epsilon$ and release output based on controllers’ privacy preferences; 2) we introduce a utility signaling protocol that helps analysts interpret the impact of DP on their downstream tasks. We implement the algorithm and the protocol inside a new platform built on top of a data escrow, which allows controllers to control dataflows while maintaining high performance. We demonstrate our contributions through an IRB-approved user study, extensive experimental evaluations, and comparison with other DP platforms. All in all, our work contributes to making DP easier to use by lowering adoption barriers.
PVLDB Reference Format:
PVLDB, 14(1): XXX-XXX, 2020.
doi:XX.XX/XXX.XX This work is licensed under the Creative Commons BY-NC-ND 4.0 International License. Visit https://creativecommons.org/licenses/by-nc-nd/4.0/ to view a copy of this license. For any use beyond those covered by this license, obtain permission by emailing info@vldb.org. Copyright is held by the owner/author(s). Publication rights licensed to the VLDB Endowment.
Proceedings of the VLDB Endowment, Vol. 14, No. 1 ISSN 2150-8097.
doi:XX.XX/XXX.XX
@@ 1. Introduction
Differential privacy (DP) (Dwork, 2006; Dwork et al., 2006b) is a promising tool to analyze data about individuals, but it remains challenging to use in practice. In a typical DP deployment, a data analyst runs queries on a dataset hosted on a platform that is accessible to a data controller. The dataset contains information about individuals, and releasing this information risks revealing the identity of those individuals. The data controller wants to protect individuals’ privacy while releasing useful output to the data analyst. DP guarantees that the output of the query is likely to be the same if we change or remove any individual’s record in the dataset. This is achieved by adding calibrated noise to the output according to a privacy loss budget $\epsilon$: smaller $\epsilon$ leads to more accurate output but weaker privacy guarantees. Despite some successful deployments of DP (Abowd, 2018; Apple, 2017; Erlingsson et al., 2014; Ding et al., 2017), a few roadblocks prevent broader DP deployment. In this paper, we consider three such roadblocks:
@@
Challenge 1: Choosing $\epsilon$. Controllers need to decide the level of privacy protection through $\epsilon$; however, it is challenging for the controller to reason about how much privacy each individual represented in the dataset receives with respect to $\epsilon$.  Dwork et al. showed that controllers have no clear consensus on how to choose $\epsilon$ or how to approach this problem in a principled manner.
@@
Our contribution. In this paper, we observe the crux of the problem is that $\epsilon$ does not directly translate into the degree of privacy loss incurred to specific individuals in the dataset. In response, we introduce a method to translate the effect of choosing $\epsilon$ into a privacy risk indicator (PRI) associated with each individual. Comparing any two individuals, the one with the higher PRI has greater privacy loss; thus, controllers learn how releasing a specific query output affects individuals’ privacy. And because PRI is connected to $\epsilon$, there is an opportunity to automatically find an $\epsilon$ that produces a profile of privacy risk indicators consistent with the controller’s privacy preferences; we exploit this opportunity by designing an algorithm, called Find-$\epsilon$-from-PRI, to choose $\epsilon$ and release DP output given a simple description of controllers’ privacy preferences.
Challenge 2: Interpreting utility of DP output. Analysts submit queries and receive noisy output that deviates from the raw output, i.e., the ground truth output if DP had not been applied. Without access to such ground truth, analysts do not know the impact of DP on the quality of their downstream tasks, which is the reason why they sought the DP output in the first place. This hinders the analysts’ ability to use the DP output in practice.
Our contribution. In this paper, we propose a utility signaling protocol to help analysts interpret the impact of DP on their downstream tasks while still allowing controllers to fully control data release. In the protocol we propose, the analyst delegates their downstream task with an associated evaluation function to the platform. With the controller’s permission, the platform invokes the evaluation function on both the DP output and the raw output and releases both results, thereby signaling the utility loss on the analyst’s downstream task due to noise introduced by DP.
Challenge 3: Deploying our solutions in a DP platform. Deploying our solutions to the previous two challenges brings a new challenge on the platform design. Specifically, implementing the new Find-$\epsilon$-from-PRI algorithm (addressing Challenge 1) and utility signaling protocol (addressing Challenge 2) calls for a platform that meets the four requirements: (1) it allows controllers to control dataflows, (2) it implements necessary security measures to protect against unintended dataflows, (3) it facilitates controllers to inspect dataflows, and (4) it adds little overhead. Various DP platforms have been proposed, but none meet all four requirements. For instance, the central DP model assumes adversaries are incapable of corrupting the platform and forcing unintended dataflows, which is unrealistic in practice (Near et al., 2021); platforms leveraging cryptographic primitives (Narayan and Haeberlen, 2012; Bater et al., 2016; Agarwal et al., 2018; Roth et al., 2019; Roy Chowdhury et al., 2020; Wang et al., 2022; Bater et al., 2018) make fewer assumptions about the adversaries, but often have significant computation overhead.
Our contribution. We propose a new platform built on top of a data escrow (Xia et al., 2022) which enables the implementation of the Find-$\epsilon$-from-PRI algorithm and utility signaling protocol, meeting the four platform requirements we curated. We show that the escrow-based platform hits a new point in the design space, offering a better trade-off between security, usability, and performance.
[…]
Summary of Contributions and Related Work
Helping controller choose $\epsilon$ using PRI. Many approaches have been proposed to help controllers choose $\epsilon$ (Hsu et al., 2014; Ge et al., 2019; Dwork et al., 2019; Lee and Clifton, 2011; Murtagh et al., 2018; Nanayakkara et al., 2022; Hay et al., 2016; Thaker et al., 2020; John et al., 2021; Ligett et al., 2017), however, none concentrate on communicating the effects of $\epsilon$ on specific individuals represented in the dataset, as we do. In addition, approaches for communicating the privacy guarantees offered by DP are mainly targeted toward end-users who share their data (Nanayakkara et al., 2023; Franzen et al., 2022), whereas our approach focuses on helping controllers, who can only make an informed choice of $\epsilon$ if they understand its privacy implications on those end-users.
Helping analyst via utility signaling protocol. Adding DP-style noise to the output affects the quality and usability of the data for the analyst’s downstream tasks (Hotz et al., 2022). For instance, the adoption of DP by the US Census has raised major concerns regarding the reliability of the census statistics after noise infusion and its effect on research and policy-making (Ruggles et al., 2019; Santos-Lozada et al., 2020; Kenny et al., 2021; ABOWD, [n.d.]). Our utility signaling protocol takes a step forward in helping analysts interpret the effect of DP on the released output.
Escrow-based DP platform. Various DP platforms have been proposed (Dwork et al., 2006a; Narayan and Haeberlen, 2012; Bater et al., 2016; Agarwal et al., 2018; Roth et al., 2019; Roy Chowdhury et al., 2020; Wang et al., 2022; Bater et al., 2018; Nguyen et al., 2020; Bittau et al., 2017), but none meets the requirements for deploying our solutions to help controllers and analysts. Our escrow-based platform meets those requirements, enabling practical deployment of DP.
@@
Outline. The rest of the paper is organized as follows. We present the new privacy risk indicator in Section 2 and the algorithm we designed to automatically find $\epsilon$ and release DP output in Section 3. Section 4 presents the utility signaling protocol and Section 5 presents the escrow-based platform. We finish with evaluation (Section 6), related work (Section 7) and conclusions in Section 8.
@@ 2. Privacy Risk Indicator (PRI)
Background. Differential privacy (DP) (Dwork, 2006; Dwork et al., 2006b) is a mathematical definition of privacy that bounds the effect a single individual has on the output of a query over a dataset of records, where each record represents an individual. Formally, a randomized mechanism $\mathcal{M}$ satisfies $\epsilon$-differential privacy if for any pair of neighboring datasets $x$ and $x^{\prime}$ where $x$ and $x^{\prime}$ differ by a single record, and for all sets of possible output $O\subseteq Range(\mathcal{M})$:
$\Pr(\mathcal{M}(x)\in O)\leq e^{\epsilon}\Pr(\mathcal{M}(x^{\prime})\in O).$
One way of achieving differential privacy is by constructing a mechanism that adds noise to the query output. The amount of noise is controlled by a tunable privacy loss budget parameter, $\epsilon$; smaller $\epsilon$ leads to larger noise and less accurate output: the choice of $\epsilon$ determines the privacy-accuracy tradeoff (Near et al., 2021).
@@ 2.1. Why Choosing $\epsilon$ is Hard and PRI Overview
Controllers decide what query output to release through $\epsilon$, which impacts individuals’ privacy. The parameter $\epsilon$ bounds the maximum privacy loss for any individual, but it is challenging for controllers to reason about the actual privacy implications of releasing an output for specific individuals in the dataset after choosing $\epsilon$ (Dwork et al., 2019).
We hypothesize that controllers can only comfortably choose $\epsilon$ if they understand how each individual’s privacy is affected by that choice. To this end, we define a privacy risk indicator (PRI), which indicates the amount of privacy loss of each individual relative to other individuals in the dataset, thus providing a direct indication of the privacy impact of releasing a query output. We construct PRI so that it is directly connected with the concept of privacy loss, which we explain next. This connection is key to both providing a principled understanding of the new indicator and designing strategies to find a suitable $\epsilon$. One strategy includes an algorithm to find $\epsilon$ based on the controller’s privacy preferences expressed in terms of PRIs, which we will explain in Section 3.
@@ 2.2. Privacy Risk Indicator: Formal Definition
We denote an $\epsilon$-DP mechanism as $\mathcal{M}:\mathcal{X}^{n}\times\mathcal{Q}\rightarrow\mathcal{Y}$, where $\mathcal{Q}$ is the query space and $\mathcal{Y}$ is the output space of $\mathcal{M}$. Specifically, we consider numeric query $q:\mathcal{X}^{n}\rightarrow\mathbb{R}^{k}$ that maps a dataset to $k$ real numbers. Given a query $q\in\mathcal{Q}$ and a query output $r\sim\mathcal{M}(x,q,\epsilon)$ of mechanism $\mathcal{M}$ which takes the dataset $x$, the query $q$ and the privacy loss budget $\epsilon$ as input, we define the privacy risk indicator of individual $i\in[n]$ as follows:
@@
$PRI_{i}=\lVert r-q(x_{-i})\rVert_{1},$
where $x_{-i}$ is the neighboring dataset of $x$ without $i$’s record $x_{i}$.
Connection between PRI and privacy loss. Next, we show that when the Laplace Mechanism (Dwork et al., 2014) is used to compute the query output $r$, PRI becomes a relative indicator of individual privacy loss. First, we show the definition of privacy loss in DP (Dwork et al., 2014).
Definition 0 (Privacy Loss).
Given two neighboring datasets $x,x^{\prime}\in\mathcal{X}^{n}$ such that $\lVert x-x^{\prime}\rVert_{1}\leq 1$, for an output $r\sim\mathcal{Y}$, the privacy loss incurred upon observing $r$ is defined as:
$\mathcal{L}_{\mathcal{M}(x,q,\epsilon)\|\mathcal{M}(x^{\prime},q,\epsilon)}^{(r)}=\ln\left(\frac{\Pr\left[\mathcal{M}(x,q,\epsilon)=r\right]}{\Pr\left[\mathcal{M}(x^{\prime},q,\epsilon)=r)\right]}\right).$
Next, we introduce the definition of Laplace Mechanism (Dwork et al., 2014), an $\epsilon$-differentially private mechanism that we use to compute the query output $r$.
Definition 0 (Laplace Mechanism).
@@
where $Z_{i}$ are i.i.d random variables drawn from the Laplace distribution with probability density function $p_{b}(z)=\frac{1}{2b}\exp(-|z|/b)$ and the scale $b=\frac{GS_{q}}{\epsilon}$  [^1] The global sensitivity of q in the $\ell_{1}$ norm is defined as $GS_{q}=\max|q(x)-q(x^{\prime})|$, where $x$ and $x^{\prime}$ are neighboring datasets in $\mathcal{X}^{n}$..
Using the Laplace Mechanism, we express the relative privacy loss of individual $i$ as:
$\mathcal{L}_{\mathcal{M_{L}}(x,q,\epsilon)\parallel\mathcal{M_{L}}(x_{-i},q,\epsilon)}^{(r)}$ $=\ln\left(\prod_{j=1}^{k}\frac{\exp(-\epsilon|r_{j}-q(x)_{j}|/GS_{q})}{\exp(-\epsilon|r_{j}-q(x_{-i})_{j}|/GS_{q})}\right)$
$=\frac{\epsilon(\lVert r-q(x_{-i})\rVert_{1}-\lVert r-q(x)\rVert_{1})}{GS_{q}}.$
@@
$\mathcal{L}_{\mathcal{M_{L}}(x,q,\epsilon)\parallel\mathcal{M_{L}}(x_{-i},q,\epsilon)}^{(r)}>\mathcal{L}_{\mathcal{M_{L}}(x,q,\epsilon)\parallel\mathcal{M_{L}}(x_{-j},q,\epsilon)}^{(r)}.$
The proof is straightforward: if $PRI_{i}>PRI_{j}$, then $\lVert r-q(x_{-i})\rVert_{1}$ $\lVert r-q(x_{-j})\rVert_{1}$, thus the inequality holds. Note that since we are using the Laplace Mechanism, which is $\epsilon$-DP, the privacy loss of any individual will always be bounded by $\epsilon$.
@@
PRI gives controllers additional information to decide what to release. The PRI becomes an additional indicator associated with each individual in the dataset that permits controllers to directly observe the effect of releasing a query output. In this paper, we show how PRI helps controllers decide which output to release.
PRI is data-dependent and thus leaks information about individuals. Controllers already have access to individuals’ raw data, so when they gain access to the PRI, they do not learn anything extra about the individuals. But this means the PRIs must remain private to the controllers and never be shared with anyone else.

### v3

# Making Differential Privacy Easier to Use for Data Controllers using a Privacy Risk Indicator

@@ Abstract.
Differential privacy (DP) enables private data analysis but is difficult to use in practice. In a typical DP deployment, data controllers manage individuals’ sensitive data and are responsible for answering data analysts’ queries while protecting individuals’ privacy; they do so by choosing $\epsilon$, the privacy loss budget, which controls how much noise to add to the query output. However, it is challenging for data controllers to choose $\epsilon$ because of the difficulty of interpreting the privacy implications of such a choice on the individuals they wish to protect.
To address this challenge, we first derive a privacy risk indicator (PRI) directly from the definition of ex-post per-instance privacy loss in the DP literature. The PRI indicates the impact of choosing $\epsilon$ on individuals’ privacy. We then leverage the PRI to design an algorithm to choose $\epsilon$ and release query output based on data controllers’ privacy preferences. We design a modification of the algorithm that allows releasing both the query output and $\epsilon$ while satisfying differential privacy, and we propose a solution that bounds the total privacy loss when using the algorithm to answer multiple queries without requiring controllers to set the total privacy loss budget. We demonstrate our contributions through an IRB-approved user study and experimental evaluations that show the PRI is useful for helping controllers choose $\epsilon$ and our algorithms are efficient. Overall, our work contributes to making DP easier to use for controllers by lowering adoption barriers.
Differential privacy, data controller, privacy loss budget
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/18/06ccs: Information systems Data management systemsccs: Security and privacy Privacy protectionsccs: Security and privacy Usability in security and privacy
@@ 1. Introduction
Differential privacy (DP) (Dwork, 2006; Dwork et al., 2006) is a promising tool to analyze data about individuals, but it remains challenging to use in practice. In a typical DP deployment, a data analyst runs queries on a dataset containing sensitive information about individuals that is accessible to a data controller. We define a data controller as an agent with the capacity to (1) directly access and run analysts’ queries on individuals’ raw data (2) make decisions on what query output to release to the analysts since the output may reveal the identity of individuals. The controller is responsible for protecting individuals’ privacy while releasing useful output to the analysts. DP guarantees that the output of the query is likely to be the same if we change or remove any individual’s record in the dataset. This is achieved by adding calibrated noise to the output according to a privacy loss budget $\epsilon$: larger $\epsilon$ leads to more accurate output but weaker privacy guarantees. Despite some successful deployments of DP (Abowd, 2018; Apple, 2017; Erlingsson et al., 2014; Ding et al., 2017), a major challenge exists which prevents broader DP deployment.
@@
Challenge of choosing $\epsilon$. Controllers need to choose the level of privacy protection through $\epsilon$; however, it is challenging for the controller to reason about how much privacy each individual represented in the dataset receives with respect to $\epsilon$.  Dwork et al. (Dwork et al., 2019) showed that controllers have no clear consensus in choosing $\epsilon$ or approaching this problem in a principled manner.
Our Contributions
@@
Designing PRI and the Find-$\epsilon$-from-PRI algorithm. In this paper, we observe the crux of the problem is that $\epsilon$ does not directly translate into the degree of privacy loss incurred to specific individuals in the dataset. We hypothesize knowing the effect of choosing $\epsilon$ on each individual’s privacy would help controllers make the choice. Therefore, we introduce a method to translate the effect of choosing $\epsilon$ into a privacy risk indicator (PRI) associated with each individual. The PRI is a relative indicator of ex-post per-instance privacy loss (Redberg and Wang, 2021), which is the realized privacy loss incurred to an individual upon releasing a specific output. Comparing any two individuals, the one with the higher PRI has greater privacy loss; thus, controllers learn how releasing a specific query output affects each individual’s privacy. And because PRI is connected to $\epsilon$, there is an opportunity to automatically find an $\epsilon$ that produces a profile of PRIs consistent with the controller’s privacy preferences; we exploit this opportunity by designing an algorithm, called Find-$\epsilon$-from-PRI, to choose $\epsilon$ and release query output given a simple description of controllers’ privacy preferences.
Privately releasing $\epsilon$. The Find-$\epsilon$-from-PRI algorithm only releases the DP output and not the chosen $\epsilon$, since $\epsilon$ is chosen based on the PRIs, which are data-dependent, and releasing $\epsilon$ directly may undermine the DP guarantees. Thus we introduce a modification of the algorithm, called Find-and-release-$\epsilon$-from-PRI, which permits releasing both the query output and $\epsilon$ while preserving DP.
Bounding privacy loss across multiple queries. In many cases, analysts submit multiple queries to the controller to obtain different information about the dataset. Besides the challenge of choosing $\epsilon$ for each query, controllers also face the challenge of setting the global privacy budget which bounds the maximum privacy loss for a dataset. We propose a solution that does not require controllers to quantify the maximum acceptable privacy loss; instead, it dynamically bounds the privacy loss after answering (or denying) each query. We construct a privacy odometer (Rogers et al., 2016) to provide an end-to-end DP guarantee for our solution.
Evaluation. In our evaluation, we test the hypothesis that the new privacy risk indicator helps controllers choose $\epsilon$ by conducting an IRB-approved user study with 56 participants. We show that our algorithms are efficient and scalable, and we conduct ablation studies to analyze the effect of varying controllers’ privacy preferences.
@@
Outline. The rest of the paper is organized as follows. We introduce the background of differential privacy and the different agents in standard DP deployment in Section 2. We present the new privacy risk indicator in Section 3 and the algorithm we designed to automatically find $\epsilon$ and release DP output in Section 4. In Section 5, we introduce the modified algorithm to privately release $\epsilon$ and our solution to dynamically bound privacy loss across multiple queries. We finish with evaluation (Section 6), related work (Section 7) and conclusions in Section 8.
2. Background
In this section, we first introduce the definition of $(\epsilon,\delta)$-differential privacy (Section 2.1), then explain the different agents in a standard DP deployment (Section 2.2).
2.1. Differential Privacy
Differential privacy (DP) (Dwork, 2006; Dwork et al., 2006) is a mathematical definition of privacy that bounds the effect a single individual has on the output of a query over a dataset of records, where each record represents an individual. Formally, for an $\epsilon,\delta\geq 0$, a randomized mechanism $\mathcal{M}$ satisfies $(\epsilon,\delta)$-differential privacy if for any pair of neighboring datasets $x$ and $x^{\prime}$ where $x$ and $x^{\prime}$ differ by a single record, and for all sets of possible output $O\subseteq Range(\mathcal{M})$:
$\Pr(\mathcal{M}(x)\in O)\leq e^{\epsilon}\Pr(\mathcal{M}(x^{\prime})\in O)+\delta.$
When the failure parameter $\delta$ is 0, $\mathcal{M}$ satisfies $\epsilon$-differential privacy; otherwise $\delta$ is usually set to be a very small probability (e.g., $\frac{1}{n^{2}}$ where $n$ is the number of records). One way of achieving DP is by constructing a mechanism that adds random noise to the query output. The amount of noise is controlled by a tunable privacy loss budget parameter, $\epsilon$. Smaller $\epsilon$ leads to more noise and less accurate output: the choice of $\epsilon$ determines the privacy-accuracy tradeoff in DP (Near et al., 2021).
There are two important properties of differential privacy, post processing and sequential composition, which are building blocks for designing sophisticated mechanisms that satisfy DP.
Post processing (Dwork et al., 2014). DP is immune to post processing, which means any computation on the output of a DP mechanism will not undermine the DP guarantees. Formally, let $\mathcal{M}(x)$ be an $\epsilon$-DP mechanism. For any arbitrary function $f$, $f(\mathcal{M}(x))$ is $\epsilon$-DP.
Sequential composition (Dwork et al., 2014). Sequentially combining multiple DP mechanisms on the same input dataset results in quantifiable degradation of privacy. Formally, given $k$ independent mechanisms $\mathcal{M}_{1},\ldots,\mathcal{M}_{k}$ that satisfy $\epsilon_{1},\ldots,\epsilon_{k}$-DP, then any arbitrary function $g(\mathcal{M}_{1},\ldots,\mathcal{M}_{k})$ is $\sum_{i=1}^{i=k}\epsilon_{i}$-DP.
2.2. Agents in DP Deployment
Figure 1. Agents in standard DP deployment
Figure 1 shows a diagram of a standard DP deployment, i.e., the central DP model (Near et al., 2021) (Wagh et al., 2021); we define a dataflow as a function that passes data from one entity to another. Individuals send their raw data to a data controller. The data analyst submits queries to the controller in order to obtain information about the individuals. The goal of the controller is to release useful output to the analyst while protecting individuals’ privacy. To achieve this, the controller answers the analyst’s queries using DP mechanisms, and they control both how much protection they want to offer to individuals and how accurate the output should be by setting the $\epsilon$ parameter when applying DP to answer each query. In addition, the controller needs to set a global privacy budget to control the maximum acceptable privacy loss incurred on the dataset; answering each query consumes some budget, once the budget is reached, the dataset is retired and can no longer be queried.
One notable example of the central DP model is the application of DP by the US Census in 2020 (Abowd, 2018). In this case, the US Census Bureau is the data controller who has access to individuals’ (i.e., the US population) data. The data analysts’ queries are pre-determined by the Census Bureau; the census data is the output of various census queries (e.g., population counts in different geographical levels by age, ethnicity, race, and sex). The Census Bureau sets the privacy loss budget $\epsilon$ and allocates the budget to each query (Bureau, 2021b)  [^1] Technically, the Census Bureau allocates the budget using the privacy parameter $\rho$ in zero-concentrated differential privacy (Bun and Steinke, 2016), which can be converted to the $\epsilon$ in $(\epsilon,\delta)$-differential privacy. We omit the details for simplicity., then uses DP to release the query output. The Census Bureau made considerable effort to choose $\epsilon$ and allocate the budget, including testing different $\epsilon$ values on a demonstration census dataset and seeking extensive feedback from experts (Bureau, 2021a). However, when deploying DP in practice, choosing $\epsilon$ is an extremely challenging task for controllers without the assistance of DP experts (Dwork et al., 2019).
While there are other DP deployment models available, i.e., the local DP (Kasiviswanathan et al., 2011) or shuffler model (Erlingsson et al., 2019), where the agents interact differently (e.g., individuals no longer send their raw data to the controller), in this paper we concentrate on addressing the controller’s challenges of choosing $\epsilon$ in the central model.
@@ 3.1. Why Choosing $\epsilon$ is Hard and PRI Overview
Controllers decide what query output to release through $\epsilon$, which impacts individuals’ privacy. The parameter $\epsilon$ bounds the maximum privacy loss for any individual a priori before running the DP mechanism to produce the query output, but it is challenging for controllers to reason about the concrete privacy implications for individuals represented in the dataset when releasing a specific output ex-post, i.e., after applying the DP mechanism.
We hypothesize that controllers can only comfortably choose $\epsilon$ if they understand how each individual’s privacy is affected by that choice. To this end, we define a privacy risk indicator (PRI), which indicates the amount of privacy loss of each individual relative to other individuals in the dataset, thus providing a direct indication of the privacy impact of releasing a query output. The PRI is an extension to a more fine-grained characterization of the privacy loss in standard DP, thus it does not depart from or replace the standard DP theory. We construct PRI so that it is directly connected with the concept of ex-post per-instance privacy loss (Redberg and Wang, 2021), which we explain next in Section 3.2. This connection is key to both providing a principled understanding of the new indicator and designing strategies to find a suitable $\epsilon$. One strategy includes an algorithm to find $\epsilon$ based on the controller’s privacy preferences expressed in terms of PRIs, which we will explain in Section 4.
@@ 3.2. Privacy Risk Indicator: Formal Definition
We denote an $(\epsilon,\delta)$-DP mechanism as $\mathcal{M}:\mathcal{X}^{n}\times\mathcal{Q}\rightarrow\mathcal{Y}$, where $\mathcal{Q}$ is the query space and $\mathcal{Y}$ is the output space of $\mathcal{M}$. Specifically, we consider numeric query $q:\mathcal{X}^{n}\rightarrow\mathbb{R}^{k}$ that maps a dataset to $k$ real numbers. Given a query $q\in\mathcal{Q}$ and a query output $r\sim\mathcal{M}(x,q,\epsilon,\delta)$ of mechanism $\mathcal{M}$ which takes the dataset $x$, the query $q$ and $(\epsilon,\delta)$ as input, we define the privacy risk indicator of an individual $i\in[n]$ as follows:
@@
$PRI_{i}=\lVert r-q(x_{-i})\rVert_{v},$
where $x_{-i}$ is the neighboring dataset of $x$ without $i$’s record $x_{i}$. The value of $v$ depends on the specific mechanism $\mathcal{M}$: if $\mathcal{M}$ is the Laplace Mechanism (Dwork et al., 2014), $v=1$ (i.e., the $\ell_{1}$ norm); if $\mathcal{M}$ is the Gaussian Mechanism (Dwork et al., 2014), $v=2$ (i.e., the $\ell_{2}$ norm).
We note that one limitation of our formulation of PRI is that it only applies to numeric queries, but not categorical queries that return an element from a set, since both $r$ and $q(x_{-i})$ must be in $\mathbb{R}^{k}$. One interesting direction is to explore a more generic formulation of PRI that adapts to both types of queries.
Connection between PRI and privacy loss. Next, we show the connection between PRI and ex-post per-instance privacy loss (Redberg and Wang, 2021), which is the privacy loss conditioned on a fixed output, an individual, and a dataset. Specifically, we show that when the Laplace Mechanism (Dwork et al., 2014) or the Gaussian Mechanism (Dwork et al., 2014) is used to compute the query output $r$, PRI becomes a relative indicator of ex-post per-instance privacy loss.
Definition 0 (Ex-post per-instance privacy loss (Redberg and Wang, 2021)).
For a randomized mechanism $\mathcal{M}$ and a query $q\in\mathcal{Q}$, the ex-post per-instance privacy loss given a realized query output $r\sim\mathcal{Y}$, a fixed dataset $x$, and an individual’s record $x_{i}\in x$ is
$\epsilon(r,x,x_{-i})=\left|\ln\left(\frac{\Pr[\mathcal{M}(x,q)=r]}{\Pr[\mathcal{M}(x_{-i},q)=r]}\right)\right|$
Note the difference between $\epsilon$, which is used as input to a DP mechanism, and $\epsilon(\cdot)$, which is a function that denotes the privacy loss conditioned on some given input, i.e., ex-post per-instance privacy loss. Intuitively, using ex-post per-instance privacy loss answers the question: what is the realized privacy loss incurred to each individual in the dataset when $r$ is released? Note that the ex-post per-instance privacy loss is a more fine-grained characterization of the privacy loss in standard DP; if $\epsilon(\cdot)\leq\epsilon$ for all $r\subseteq Range(\mathcal{M}),x\in\mathcal{X}^{n}$, and $x$’s neighboring dataset $x^{\prime}$ where $\lVert x-x^{\prime}\rVert_{1}\leq 1$, then $\mathcal{M}$ satisfies $\epsilon$-DP (similar argument can be made for $(\epsilon,\delta)$-DP). We derive our notion of PRI from the ex-post per-instance privacy loss so that it describes the realized individual privacy loss when an output of a DP mechanism is released. The PRI works as a relative indicator of privacy loss for both Laplace Mechanism and Gaussian Mechanism, which we explain next.
Definition 0 (Laplace Mechanism (Dwork et al., 2014)).
@@
where $Z_{i}$ are i.i.d random variables drawn from the Laplace distribution with probability density function $p_{b}(z)=\frac{1}{2b}\exp(-|z|/b)$ and the scale $b=\frac{\Delta_{1}}{\epsilon}$  [^2] The global sensitivity of q in the $\ell_{1}$ norm is defined as $\Delta_{1}=\max\lVert q(x)-q(x^{\prime})\rVert_{1}$, where $x$ and $x^{\prime}$ are neighboring datasets in $\mathcal{X}^{n}$.. The Laplace Mechanism satisfies $\epsilon$-DP.
When $r$ is released using the Laplace Mechanism, the ex-post per-instance privacy loss is computed as:
$\epsilon_{L}(r,x,x_{-i})$ $=\left|\ln\left(\prod_{j=1}^{k}\frac{\exp(-\epsilon|r_{j}-q(x)_{j}|/\Delta_{1})}{\exp(-\epsilon|r_{j}-q(x_{-i})_{j}|/\Delta_{1})}\right)\right|$
$=\left|\epsilon\frac{\lVert r-q(x_{-i})\rVert_{1}-\lVert r-q(x)\rVert_{1}}{\Delta_{1}}\right|.$
@@
$\epsilon_{L}(r,x,x_{-i})>\epsilon_{L}(r,x,x_{-j})$
The proof is straightforward: $PRI_{i}>PRI_{j}$ means $\lVert r-q(x_{-i})\rVert_{1}$ $\lVert r-q(x_{-j})\rVert_{1}$ (the PRI is in $\ell_{1}$ norm since Laplace Mechanism is used), thus the inequality holds.
Definition 0 (Gaussian Mechanism (Dwork et al., 2014)).
Given a query $q:\mathcal{X}^{n}\rightarrow\mathbb{R}^{k}$, dataset $x\in\mathcal{X}^{n}$ and $\epsilon,\delta>0$, the Gaussian Mechanism is defined as:
$\mathcal{M}_{G}(x,q,\epsilon,\delta)=q(x)+(Z_{1},\ldots,Z_{k}),$
where $Z_{i}$ are i.i.d random variables drawn from the Gaussian distribution $\mathcal{N}(\sigma^{2})$ with center 00 and variance $\sigma^{2}=\frac{2\Delta_{2}^{2}\ln(1.25/\delta)}{\epsilon^{2}}$  [^3] The global sensitivity of q in the $\ell_{2}$ norm is defined as $\Delta_{2}=\max\lVert q(x)-q(x^{\prime})\rVert_{2}$, where $x$ and $x^{\prime}$ are neighboring datasets in $\mathcal{X}^{n}$.. The Gaussian Mechanism satisfies $(\epsilon,\delta)$-DP.
When $r$ is released using the Gaussian Mechanism, the ex-post per-instance privacy loss is computed as:
$\epsilon_{G}(r,x,x_{-i})$ $=\left|\ln\left(\frac{\exp(-\lVert r-q(x)\rVert_{2}^{2}/2\sigma^{2})}{\exp(-\lVert r-q(x_{-i})\rVert_{2}^{2}/2\sigma^{2})}\right)\right|$
$=\left|\frac{\lVert r-q(x_{-i}\rVert_{2}^{2}-\lVert r-q(x)\rVert_{2}^{2}}{2\sigma^{2}}\right|.$
Claim: For all individuals $i,j$ where $i\neq j$, if $PRI_{i}>PRI_{j}$, then
$\epsilon_{G}(r,x,x_{-i})>\epsilon_{G}(r,x,x_{-j})$
The proof is similar to that of the Laplace Mechanism, except now the PRI is in $\ell_{2}$ norm. Hence $PRI_{i}>PRI_{j}$ means $\lVert r-q(x_{-i})\rVert_{2}$ $\lVert r-q(x_{-j})\rVert_{2}$.
@@
PRI gives controllers additional information to decide what to release. The PRI becomes an additional indicator associated with each individual in the dataset that permits controllers to directly observe the effect of releasing a query output. The PRI is private information available to controllers only. In this paper, we show how PRI helps controllers decide which output to release.
PRI is data-dependent and thus leaks information about individuals. Controllers already have access to individuals’ raw data, so when they gain access to the PRIs, they do not learn anything extra about the individuals. Similar to how controllers would not release the individuals’ raw data, they will not release the PRIs. PRIs must remain private to the controllers and never be shared with any other agents.